OpenStudy (anonymous):

The equation x2 + y2 − 2x + 2y − 1 = 0 is the general form of the equation of a circle. What is the standard form of the equation?

OpenStudy (anonymous):

anwsers (x − 1)2 − (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 3 (x + 1)2 + (y − 1)2 = 1 (x − 1)2 + (y + 1)2 = 1

OpenStudy (mathmate):

The standard form of a circle is (x-h)^2+(y-k)^2=r^2 Here h=1, k=-1, and r^2=3 Can you figure out the choice?

OpenStudy (anonymous):

im so confused right now